solve and check your answer
b/4+b/3=7

Answer by checkley77(12844)

b/4+b/3=7
(3b+4b)/4*3=7
7b/12=7
7b=7*12
b=12 ans.
Proof:
12/3+12/4=7
4+3=7
7=7